OpenStudy (crashonce):

A man walks from A to B at x m/s and back at y m/s. What is his average speed for the whole journey

hartnn (hartnn):

average speed = total distance from A to B and back/ time taken to do so

OpenStudy (crashonce):

can u give me an equation

hartnn (hartnn):

Let distance between A to B = D so total distance = numerator = D right ?

hartnn (hartnn):

numerator = 2D **

OpenStudy (crashonce):

2D/(d/x+d/y)

hartnn (hartnn):

speed = distance / time x = D/time1 time 1 = D/x time 2 = D/y

OpenStudy (crashonce):

so the final answer is 2xy/x+y?

hartnn (hartnn):

yes :)

OpenStudy (crashonce):

I need u to help me with a few more @hartnn ill tag u

hartnn (hartnn):

ok :)